The Au-Au run is close to entering its last two weeks following this weekend. The last several weks even though the uptime has been only ~40 hours per week has been very productive for us, and brahms has collected lots of good data spanning many rapidities. f The overall recent progress can be followed on the http://pii3.brahms.bnl.gov/~operator/shifts/RunStatistics2001.htm page that roughly reflect the near-term run-plan as well as data collected in the last several weeks. This plan is updated daily for the shifters but not necessarely so for the web page. During the open maintnance yesterday a few activities were carried out. a) The MRS platform i.. TPM1-D5-Tpm2 which rest on a seperately moveable carriage was move backward by ~50 cm (+- .5 cm) mainly to allow us to go to 30 deg. as well as reduce the rather high occupancy at 40-45 deg. This could lead to a measurement of HBT at 40 deg. c) The attempted repair of the T5-II module was unsuccesfull. It is not possible in such short time, and without the experts to carry this out. This unfortunately means the Back Front Spectrometer is not fully op. (T3 is already marginal effeciencient). Over this weekend if beam is available at normal duty factor we should be able to complete the 'shorter' survey measurement and can start longer runs to collect high statistics next week. Preparations for the pp running is continuing. A start counter to be placed in fron of D1 at approx ~ 2.3 m a2 and covering (most) of the D1 entrance is underway. The frame is in the shops, and together with the H1 and H2 slats hopefully a good track trigger can be devised. Please be advised Monday is Memorial Day an official lab holiday. ------------------------------------------------------ Flemming Videbaek Physics Department Brookhaven National Laboratory tlf: 631-344-4106 fax 631-344-1334 e-mail: videbaek@bnl.gov
This archive was generated by hypermail 2b30 : Fri Nov 09 2001 - 21:09:17 EST